# Rigidity of Gradient Shrinking Ricci Solitons with Bach-like Flatness and Related Variational Formulas

### v2

@@ 10 Variational Origins of U and V
The classical Bach tensor arises naturally in a number of ways. One approach is to start with the conformal Einstein equations and then study integrability conditions via divergence operations. A necessary condition to be conformally Einstein is for the Bach tensor $\mathrm{B}_{ij}$ to vanish. The approach taken by R. Bach’s [bach:1921] was to consider the first variation of the quadratic Riemannian functional
@@
so the Bach tensor is a critical point of a Riemannian functional whose energy depends on how much a metric deviates from being conformally flat. In this section, we identify Bach-like tensors as the Euler–Lagrange gradients of natural quadratic curvature functionals. We will work in the standard transverse–traceless (TT) gauge for the second variation, imposing for a variation $h$
@@
It is worth recalling the Lichnerowicz Laplacian as it will appear in our second variation formulas. This operator $\Delta_{L}$ acts on symmetric 2-tensors via the rule
$(\Delta_{L}\,h)_{ij}=(\nabla^{*}\nabla\,h)_{ij}+2\,\mathrm{R}_{ikjl}\,h_{kl}-\mathrm{R}_{ik}\,h_{kj}-\mathrm{R}_{jk}\,h_{ki}.$
Here $\nabla^{*}$ denotes the divergence operator or formal adjoint of the covariant derivative $\nabla$, and $\nabla^{*}\nabla$ is the rough Laplacian. Note that sign conventions for the Laplacian vary, but we will take the convention that
$\nabla^{*}\nabla\,h=-\nabla^{k}\nabla_{k}\,h.$
@@
Then the $L^{2}$–gradient satisfies
$\mathrm{grad}\,\mathcal{F}_{\alpha,\beta}\;=\;\alpha\,\mathrm{U}\;+\;\beta\,\mathrm{V}.$
In particular, the Euler–Lagrange equation for $\mathcal{F}_{\alpha,\beta}$ is equivalent to the vanishing of the Bach-like tensor $\mathfrak{B}_{ij}=\alpha\,\mathrm{U}_{ij}+\beta\,\mathrm{V}_{ij}=0$.
@@
These standard variational identities for quadratic functionals are well-established [e.g., [besse:1987]]. Using $\mathrm{B}=\tfrac{1}{2}\,\mathrm{U}+\tfrac{1}{6}\,\mathrm{V}$,
$\delta\mathcal{F}_{\alpha,\beta}=\int\Big\langle\alpha\!\left(\tfrac{1}{2}\mathrm{U}+\tfrac{1}{6}\mathrm{V}\right)\;-\;\tfrac{\alpha/3-\beta}{2}\cdot 2\mathrm{V},\,\delta g\Big\rangle\,\mathrm{d}V_{g}=\int\langle\alpha\mathrm{U}+\beta\mathrm{V},\,\delta g\rangle\,\mathrm{d}V_{g}.$
Hence $\mathrm{grad}\,\mathcal{F}_{\alpha,\beta}=\alpha\,\mathrm{U}+\beta\,\mathrm{V}$. ∎
@@
Proposition 10.3 (Second variation of $\int R^{2}$ in TT gauge).
Let $(M^{4},g)$ be closed. For a variation $g_{t}=g+t\,h+o(t)$ with
$\mathrm{div}h=0,\qquad\mathrm{tr}_{g}h=0\quad\text{\emph{(TT gauge)}},$
the second variation of $F(g):=\int_{M}\mathrm{R}^{2}\,\,\mathrm{d}V_{g}$ is
@@
Write the first variation as $\delta F=-2\int\langle\mathrm{V},h\rangle\,\mathrm{d}V_{g}$. Then by differentiating once more, and using $\mathrm{div}h=0$, $\mathrm{tr}_{g}h=0$, the Hessian is
@@
We have the following standard TT-gauge variations:
@@
$\delta(\nabla_{i}\nabla_{j}\mathrm{R})$ $=\nabla_{i}\nabla_{j}(\delta\mathrm{R})-\tfrac{1}{2}(\nabla_{i}h_{j}{}^{k}+\nabla_{j}h_{i}{}^{k}-\nabla^{k}h_{ij})\,\nabla_{k}\mathrm{R},$
@@
$\delta\mathrm{Rc}$ $=\tfrac{1}{2}(\nabla^{*}\nabla h)_{ij}+(\mathrm{Rm}\!*h)_{ij}-(\mathrm{Rc}\circ h+h\circ\mathrm{Rc})_{ij}.$
Using these, the Hessian can be written as
$\delta^{2}F[h,h]=\int_{M}\Big(-\mathrm{R}\,\langle\Delta_{L}h,h\rangle+2\langle\mathrm{Rc},h\rangle^{2}+(\frac{1}{2}\mathrm{R}^{2}-2\Delta\mathrm{R})|h|^{2}\Big)\,\,\mathrm{d}V_{g}.$
We expand the Lichnerowicz laplacian as $\Delta_{L}h=\nabla^{*}\nabla h+2\mathrm{Rm}\!*h-\mathrm{Rc}\circ h-h\circ\mathrm{Rc}$. Integration by parts yields the result. ∎
@@
On an Einstein $4$-manifold $(M,g_{E})$ with $\mathrm{Rc}=\lambda g_{E}$ and scalar curvature $R=4\lambda$, the second variation of $\mathcal{F}_{\alpha,\beta}(g)$ in the TT-gauge is
@@
In TT gauge at an Einstein metric, the standard linearizations give
@@
Since
$\delta^{2}\!\int_{M}|\mathrm{W}|^{2}\,\mathrm{d}V_{g}=-2\int_{M}\!\langle\mathrm{B}^{\prime}(h),h\rangle\,\mathrm{d}V_{g}\quad\mathrm{and}\quad\delta^{2}\!\int_{M}\mathrm{R}^{2}\,\mathrm{d}V_{g}=-2\int_{M}\!\langle\mathrm{V}^{\prime}(h),h\rangle\,\mathrm{d}V_{g},$
@@
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=-2\alpha\!\int_{M}\langle\mathrm{B}^{\prime}(h),h\rangle\,\mathrm{d}V_{g}\;-\;(\alpha/3-\beta)\!\int_{M}\langle\mathrm{V}^{\prime}(h),h\rangle\,\mathrm{d}V_{g}.$
@@
On the Bach line $\beta=\alpha/3$, one has
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\alpha\int_{M}\left\langle\big(\Delta_{L}-\tfrac{\mathrm{R}}{3}\big)\big(\Delta_{L}-\tfrac{\mathrm{R}}{6}\big)h,\;h\right\rangle\,\mathrm{d}V_{g},$
i.e. the Hessian factorizes into two shifted Lichnerowicz factors.
@@
Let $(M^{4},g_{E})$ be Einstein with TT–spectrum of $\Delta_{L}$ contained in $\mu_{0},\infty)$. If $p_{\alpha,\beta}(\mu)>0$ for every TT–eigenvalue $\mu$ (equivalently, $\inf_{\mu\geq\mu_{0}}p_{\alpha,\beta}(\mu)>0$), then $g_{E}$ is a strict local minimizer of $\mathcal{F}_{\alpha,\beta}$ modulo diffeomorphisms and scaling. In particular, there exists a neighborhood of $g_{E}$ in which the equation $\alpha U+\beta V=0$ has no nontrivial solutions other than metrics isometric to $g_{E}$.
@@
If $R=0=W$ (Euclidean space, or a flat compact quotient), then $B^{\prime}(h)=\tfrac{1}{2}\Delta_{L}^{2}h$ and $V^{\prime}(h)\equiv 0$ on TT. Hence
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\alpha\int_{M}\langle\Delta_{L}^{2}h,h\rangle\,\,\mathrm{d}V=\alpha\int_{M}|\Delta h|^{2}\,\,\mathrm{d}V\quad(\text{TT}),$
which is strictly positive for $\alpha>0$ on compactly supported (or suitably decaying/weighted) variations.

### v3

@@ 9 Variational Origins of U and V
The classical Bach tensor arises naturally in a number of ways. One approach is to start with the conformal Einstein equations and then study integrability conditions via divergence operations. A necessary condition to be conformally Einstein is for the Bach tensor $\mathrm{B}_{ij}$ to vanish. The approach taken by R. Bach [2] was to consider the first variation of the quadratic Riemannian functional
@@
so the Bach tensor is the gradient of a Riemannian functional whose energy depends on how much a metric deviates from being conformally flat. In this section, we identify Bach-like tensors as the Euler–Lagrange gradients of natural quadratic curvature functionals. We work in the standard transverse-traceless (TT) gauge for the second variation, imposing for a variation $h$:
@@
Recall the Lichnerowicz Laplacian acting on symmetric $2$-tensors:
$(\Delta_{L}\,h)_{ij}=(\nabla^{*}\nabla\,h)_{ij}+2\,\mathrm{R}_{ikjl}\,h_{kl}-\mathrm{R}_{ik}\,h_{kj}-\mathrm{R}_{jk}\,h_{ki},$
where $\nabla^{*}\nabla\,h=-\nabla^{k}\nabla_{k}\,h$.
@@
Then the $L^{2}$-gradient satisfies $\mathrm{grad}\,\mathcal{F}_{\alpha,\beta}=\alpha\,\mathrm{U}+\beta\,\mathrm{V}$. In particular, the Euler–Lagrange equation for $\mathcal{F}_{\alpha,\beta}$ is $\mathfrak{B}_{ij}=\alpha\,\mathrm{U}_{ij}+\beta\,\mathrm{V}_{ij}=0$.
@@
Using $\mathrm{B}=\tfrac{1}{2}\,\mathrm{U}+\tfrac{1}{6}\,\mathrm{V}$:
$\delta\mathcal{F}_{\alpha,\beta}=\int\Big\langle\alpha\!\left(\tfrac{1}{2}\mathrm{U}+\tfrac{1}{6}\mathrm{V}\right)\;-\;\tfrac{\alpha/3-\beta}{2}\cdot 2\mathrm{V},\,\delta g\Big\rangle\,\mathrm{d}V_{g}=\int\langle\alpha\mathrm{U}+\beta\mathrm{V},\,\delta g\rangle\,\mathrm{d}V_{g}.\qed$
@@
Proposition 9.3 (Second variation of $\int\mathrm{R}^{2}$ in TT gauge).
Let $(M^{4},g)$ be closed. For a TT variation $g_{t}=g+th+o(t)$, the second variation of $F(g):=\int_{M}\mathrm{R}^{2}\,\,\mathrm{d}V_{g}$ is
@@
Write the first variation as $\delta F=-2\int\langle\mathrm{V},h\rangle\,\mathrm{d}V_{g}$. Differentiating once more and using TT gauge:
@@
Expanding using the standard TT-gauge variations
@@
$\delta\mathrm{Rc}$ $=\tfrac{1}{2}(\nabla^{*}\nabla h)_{ij}+(\mathrm{Rm}\!*h)_{ij}-(\mathrm{Rc}\circ h+h\circ\mathrm{Rc})_{ij},$
and integrating by parts yields the result. ∎
@@
On an Einstein $4$-manifold $(M,g_{E})$ with $\mathrm{Rc}=\lambda g_{E}$ and $\mathrm{R}=4\lambda$, the second variation of $\mathcal{F}_{\alpha,\beta}(g)$ in TT gauge is
@@
In TT gauge at an Einstein metric:
@@
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=-2\alpha\!\int_{M}\langle\mathrm{B}^{\prime}(h),h\rangle\,\mathrm{d}V_{g}\;-\;(\alpha/3-\beta)\!\int_{M}\langle\mathrm{V}^{\prime}(h),h\rangle\,\mathrm{d}V_{g}.\qed$
@@
On the Bach line $\beta=\alpha/3$:
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\alpha\int_{M}\left\langle\big(\Delta_{L}-\tfrac{\mathrm{R}}{3}\big)\big(\Delta_{L}-\tfrac{\mathrm{R}}{6}\big)h,\;h\right\rangle\,\mathrm{d}V_{g}.$
@@
Let $(M^{4},g_{E})$ be Einstein with TT-spectrum of $\Delta_{L}$ contained in $\mu_{0},\infty)$. If $p_{\alpha,\beta}(\mu)>0$ for every TT-eigenvalue $\mu$, then $g_{E}$ is a strict local minimizer of $\mathcal{F}_{\alpha,\beta}$ modulo diffeomorphisms and scaling. In particular, there exists a neighborhood of $g_{E}$ in which $\alpha\,\mathrm{U}+\beta\,\mathrm{V}=0$ has no nontrivial solutions other than metrics isometric to $g_{E}$.
@@
If $\mathrm{R}=0=\mathrm{W}$, then $\mathrm{B}^{\prime}(h)=\tfrac{1}{2}\Delta_{L}^{2}h$ and $\mathrm{V}^{\prime}(h)\equiv 0$ on TT. Hence
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\alpha\int_{M}|\Delta h|^{2}\,\,\mathrm{d}V,$
which is strictly positive for $\alpha>0$ on compactly supported variations.